GlueX TOF Meeting, June 6, 2014
GlueX Time-of-Flight Meeting
Friday, June 6, 2014
10:00 am EDT
JLab: CEBAF Center, L207
- TOF Trigger (Sascha S.) https://halldweb.jlab.org/level-1/algorithm/tof_cluster.pptx
- Review of meeting on May 22
- Installation Update (Hussein)
- checking the PMTs
- Software Discussion (all)
- What are the pressing issues?
- What are the pressing issues?
- To join via Polycom room system go to the IP Address: 184.108.40.206 (bjn.vc) and enter the meeting ID: 193936155
- To join via a Web Browser, go to the page  https://bluejeans.com/193936155
- To join via phone, use one of the following numbers and the Conference ID: 193936155
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- More information on connecting to bluejeans is available.
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2014-2Q on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2014-2Q/ .
- FSU: Paul Eugenio, Sasha Ostrovidov
- JLab: Hussein Al Ghoul, Eugene Chudakov, Mark Ito (chair), Sascha Somov, Beni Zihlmann
Review of Minutes from May 22
We looked at the minutes from May 22. Hussein told us that the HV GUI now reports when the voltages are ramping up or ramping down.
Sascha S. brought us up to date on the status of a TOF trigger.
- For bring the system up with cosmics, the closest detector is the FCAL, but an FCAL trigger will likely not help much. A trigger with external scintillator counters will likely not have good coverage of the TOF. We would like to have a TOF self-trigger of some sort.
- A total energy trigger, like that used for the FCAL is not hard to implement. That will be the first step.
- Sascha showed a slide describing a TOF trigger suitable for calibration triggers when we have beam. A 32-bit word is output from the CTP, 16-bits per layer. Each bit indicates whether there was and end-to-end coincidence in any of a group of counters.
- We discussed getting the read-out going. There are three TOF DAQ crates installed. The flashes have been read-out, but not all three crates together. Sascha S. will have time to turn to this effort next week.
- Sasha O. asked about scalers. Sascha S. described our two options
- scalers in the FADC, part of the read-out
- scalers in the discriminator, available via EPICS
- We need to ask Hovanes Egiyan about the status of these scalers.
Paul and Sasha plan to travel to JLab to work on the DAQ, most probably the week after next.
Hussein has been looking at PMTs with the scope, both ends at a time, triggering on one and looking at the other. He set the maximum voltages (hard and soft) at 1800 V and the set voltage at 1200 V. No light leaks have been found so far. He has found some cabling anomalies. He gave the update.
I started checking the modules for possible light leaks and cabling problems. Below is an image of the tested modules so far. All modules in the horizontal plane have been tested and problems were resolved. Keep in mind that TOF-N and TOF-S correspond to modules in the horizontal plane (Y), TOF-UP and TOF-DW correspond to modules in the vertical plane.
We tried to identify issues in the TOF software.
- Beni noted that the black plastic covering the FCAL and the Lexan sheet used for light pulsing of the FCAL are not in the geometry.
- Beni also noted that the iron frames are not in the geometry. These should be added.
- Beni thought that the process of taking raw EVIO data and converting it to physical quantities likely needs work. We need to talk to David to get the current status.
- Sasha O. pointed out that the calibration software still needs to be developed.